# Annabelle Michael Carrell    Albert Gong    Abhishek Shetty    Raaz Dwivedi    Lester Mackey

### v5

@@ 1 Introduction
This work is about thinning, finding a small set of representative points to accurately summarize a larger dataset. State-of-the-art thinning techniques provably improve upon uniform subsampling but only for restricted classes of kernel-based quality measures and with pessimistic dependence on the data dimension (see, e.g., Harvey & Samadi, 2014; Phillips & Tai, 2020; Alweiss et al., 2021; Dwivedi & Mackey, 2024, 2022; Shetty et al., 2022; Li et al., 2024). We introduce a new analysis for sub-Gaussian thinning algorithms that applies to any kernel and shows that one can efficiently identify a better-than-uniform set of representative points whenever the kernel or data matrix is nearly low-rank. This opens the door to a variety of impactful applications including approximate dot-product attention in transformers, accelerated stochastic gradient training, and distinguishing distributions with deep kernels in near-linear time.
@@ 2 Sub-Gaussian Thinning
Consider a fixed collection of $n_{\textup{in}}$ input points $\mathcal{X}_{\textup{in}}$ belonging to a potentially larger universe of datapoints $\mathcal{X}\triangleq\{\boldsymbol{x}_{1},\dots,\boldsymbol{x}_{n}\}$. The aim of a thinning algorithm is to select $n_{\textup{out}}$ points from $\mathcal{X}_{\textup{in}}$ that together accurately summarize $\mathcal{X}_{\textup{in}}$. This is formalized by the following definition.
@@
A thinning algorithm Alg takes as input $\mathcal{X}_{\textup{in}}$ and returns a possibly random subset $\mathcal{X}_{\textup{out}}$ of size $n_{\textup{out}}$. We denote the input and output empirical distributions by $\mathbb{P}_{\textup{in}}\triangleq\frac{1}{n_{\textup{in}}}\sum_{\boldsymbol{x}\in\mathcal{X}_{\textup{in}}}\boldsymbol{\delta}_{\boldsymbol{x}}$ and $\mathbb{P}_{\textup{out}}\triangleq\frac{1}{n_{\textup{out}}}\sum_{\boldsymbol{x}\in\mathcal{X}_{\textup{out}}}\boldsymbol{\delta}_{\boldsymbol{x}}$ and define the induced probability vectors $\boldsymbol{p}_{\textup{in}},\boldsymbol{p}_{\textup{out}}\in\boldsymbol{\Delta}_{n-1}$ over the indices $[n]$ by
@@
When $\mathcal{X}\subset{}^{d}$, we use $\mathbf{X}\triangleq[\boldsymbol{x}_{1},\ldots,\boldsymbol{x}_{n}]^{\top}\in{}^{n\times d}$ to denote the input point matrix so that
@@
Given two distributions $\mu,\nu$ and a reproducing kernel $\mathbf{k}$ (Steinwart & Christmann, 2008, Def. 4.18), the associated kernel maximum mean discrepancy (MMD) is the worst-case difference in means for functions in the unit ball $\mathbb{B}_{\mathbf{k}}$ of the associated reproducing kernel Hilbert space:
$\operatorname{MMD}_{\mathbf{k}}(\nu,\mu)$ $\triangleq\sup_{f\in\mathbb{B}_{\mathbf{k}}}|\mathbb{E}_{\boldsymbol{x}\sim\mu}f(\boldsymbol{x})-\mathbb{E}_{\boldsymbol{x}\sim\nu}f(\boldsymbol{x})|.$ (3)
When $\mu=\mathbb{P}_{\textup{in}}$ and $\nu=\mathbb{P}_{\textup{out}}$ as in Def. 1 and $\mathbf{K}\triangleq(\mathbf{k}(\boldsymbol{x}_{i},\boldsymbol{x}_{j}))_{i,j=1}^{n}\in{}^{n\times n}$ denotes the induced kernel matrix, then the MMD can be expressed as a Mahalanobis distance between $\boldsymbol{p}_{\textup{in}}$ and $\boldsymbol{p}_{\textup{out}}$:
@@
We write $\textsc{Alg}\in\mathcal{G}_{\nu,\delta}(\mathbf{K})$ and say Alg is $(\mathbf{K},\nu,\delta)$-sub-Gaussian, if Alg is a thinning algorithm, $\mathbf{K}$ is a symmetric positive semidefinite (SPSD) matrix, $\nu>0$, $\delta\in[0,1)$, and there exists an event $\mathcal{E}$ with probability at least $21-\delta/$ 2 such that, the input and output probability vectors satisfy
@@ 2.1 Examples of sub-Gaussian thinning algorithms
Perhaps the simplest sub-Gaussian thinning algorithm is uniform subsampling: by Prop. B.1, selecting $n_{\textup{out}}$ points from $\mathcal{X}_{\textup{in}}$ uniformly at random (without replacement) is $(\mathbf{K},\nu,0)$-sub-Gaussian with $\nu={\sqrt{\|{\mathbf{K}}\|_{\max}}/\sqrt{n_{\textup{out}}}}$. Unfortunately, uniform subsampling suffers from relatively poor summarization quality. As we prove in Sec. B.1.1, its root-mean-squared MMD and KMS are both $\Omega(1/\sqrt{n_{\textup{out}}})$ meaning that $n_{\textup{out}}=10000$ points are needed to achieve $1\%$ relative error.
@@
Fortunately, uniform subsampling is not the only sub-Gaussian thinning algorithm available. For example, the Kernel Halving (KH$(\delta)$) algorithm of Dwivedi & Mackey (2024) provides a substantially smaller sub-Gaussian parameter, $\nu=O({\sqrt{\log(n_{\textup{out}}/\delta)}}{/n_{\textup{out}}})$, at the cost of $n_{\textup{in}}^{2}$ runtime, while the KH-Compress$(\delta)$ algorithm of Shetty et al. (2022, Ex. 3) delivers $\nu=O({\sqrt{\log(n_{\textup{out}})\log(n_{\textup{out}}/\delta)}}{/n_{\textup{out}}})$ in only $n_{\textup{out}}^{2}$ time. We derive simplified versions of these algorithms with identical sub-Gaussian constants in Secs. B.2 and B.5 and a linear-kernel variant (LKH$(\delta)$) with $n_{\textup{in}}d$ runtime in Sec. B.3. To round out our set of examples, we show in Sec. B.6.1 that two new thinning algorithms based on the Gram-Schmidt walk of Bansal et al. (2018) yield even smaller $\nu$ at the cost of increased runtime. We call these algorithms Gram-Schmidt Thinning (GS-Thin) and GS-Compress.
@@ 3 Low-rank Sub-Gaussian Thinning
Let us unpack the three components of this result. First, Thm. 1 provides a high-probability $O(\nu\sqrt{\log(|\mathcal{I}|)})$ bound 13 on the KMS for any kernel and any sub-Gaussian thinning algorithm on any space. In particular, the non-uniform algorithms of Tab. 1 all enjoy $O(\log(n_{\textup{out}})\sqrt{\log(|\mathcal{I}|)}/n_{\textup{out}})$ KMS, a significant improvement over the $\Omega(1/\sqrt{n_{\textup{out}}})$ KMS of uniform subsampling. Second, Thm. 1 provides a refined $O(\nu\sqrt{\mathrm{rank}({\mathbf{X}_{\mathcal{I}}})\log(R_{\mathcal{I}}L_{\mathbf{K}})})$ bound 15 on KMS for datapoints in $\mathbb{R}^{d}$. For bounded data, this trades an explicit dependence on the number of query points $|\mathcal{I}|$ for a rank factor that is never larger (and sometimes significantly smaller) than $d$. We will make use of these results when approximating dot-product attention in Sec. 4.
Finally, Thm. 1 provides an $O(\nu\sqrt{r}+\sqrt{\lambda_{r+1}/n_{\textup{out}}})$ high-probability bound on kernel MMD, where the approximate rank parameter $r$ can be freely optimized. When $\mathbf{K}=(\mathbf{k}(\boldsymbol{x}_{i},\boldsymbol{x}_{j}))_{i,j=1}^{n}$ is generated by a finite-rank kernel $\mathbf{k}$, like a linear kernel $\langle{\boldsymbol{x}_{i}},{\boldsymbol{x}_{j}}\rangle$, a polynomial kernel $(1+\langle{\boldsymbol{x}_{i}},{\boldsymbol{x}_{j}}\rangle)^{p}$, or a random Fourier feature kernel (Rahimi & Recht, 2007), this guarantee becomes $O(\nu)$ and improves upon uniform subsampling whenever $\nu=o(1/\sqrt{n_{\textup{out}}})$. In this case, the non-uniform algorithms of Tab. 1 all enjoy $O(\log(n_{\textup{out}})/n_{\textup{out}})$ MMD, a significant improvement over the $\Omega(1/\sqrt{n_{\textup{out}}})$ MMD of uniform subsampling. We will revisit this finite-rank setting when studying stochastic gradient acceleration strategies in Sec. 5.
@@
Altschuler et al. (2019, Thm. 4) additionally showed that
@@ 4 Approximating Attention
Dot-product attention lies at the heart of the Transformer neural network architecture that has revolutionized natural language processing, computer vision, and speech recognition over the last decade (Vaswani et al., 2017; Dosovitskiy et al., 2021; Dong et al., 2018). Given a collection of query, key, and value vectors $(\boldsymbol{q}_{i},\boldsymbol{k}_{i},\boldsymbol{v}_{i})_{i=1}^{n}$ each in $\mathbb{R}^{d}$, dot-product attention computes the softmax matrix
@@
$\text{for}\ \mathbf{A}_{ij}\triangleq\exp(\frac{\langle{\boldsymbol{q}_{i}},{\boldsymbol{k}_{j}}\rangle}{\sqrt{d}}),\mathbf{D}=\mathop{\mathrm{missing}}{diag}(\mathbf{A}\mathbf{1}_{n}),\text{ and }\mathbf{V}_{ij}\triangleq\boldsymbol{v}_{ij}.$ (23)
@@ 4.1 Thinning attention in theory
Alg. 1 summarizes our new Thinformer module. At its heart is a new key-value attention kernel $\mathbf{k}_{\mathrm{att}}$ that mimics the special structure of the softmax matrix $\mathbf{T}$. Alg. 1 uses the attention kernel and a high-quality thinning algorithm, KH-Compress$(0.5)$, to subselect key-value pairs and then computes exact attention 22 for the key-value subset. In total, this requires only $O(d\,n_{\textup{out}}^{2})$ time to run KH-Compress$(0.5)$ and $O(d\,n\,n_{\textup{out}})$ time to compute Attention with $n$ queries and $n_{\textup{out}}$ key-value pairs. In contrast, computing the exact softmax matrix $\mathbf{T}$ with standard matrix multiplication requires $\Theta(d\,n^{2})$ time. Our next result, proved in App. F, shows that Alg. 1 also admits a strong quality guarantee for approximating $\mathbf{T}$.
@@
Table 3: Quality of T2T-ViT attention approximations on ImageNet. We report mean Top-$1$ accuracy ±$\pm 1$ standard deviation across five random seeds and mean forward pass runtime ±$\pm 1$ standard deviation across $50$ batches of $64$ images.
@@
$\|{\mathbf{\widehat{T}}-\mathbf{T}}\|_{\max}\leq\!\frac{c\exp(\frac{2R^{2}}{\sqrt{d}})\|{\mathbf{V}}\|_{2,\infty}\sqrt{\log_{2}(n_{\textup{out}})\log({12n_{\textup{out}}\log_{2}\frac{n_{\textup{in}}}{n_{\textup{out}}}})}}{n_{\textup{out}}}$ (24)
@@
the KDEformer approximation $\mathbf{\widehat{T}}_{\textrm{kde}}$ (Zandieh et al., 2023, Cor. 3.6) with $\tau=0.173+o(1)$, the HyperAttention approximation $\mathbf{\widehat{T}}_{\textrm{hyp}}$ (Han et al., 2024, Thm. 1) with no masking, and the Thinformer approximation $\mathbf{\widehat{T}}_{\textrm{thin}}$ guarantee
$\|{\mathbf{\widehat{T}}_{\textrm{kde}}-\mathbf{T}}\|_{\max}$ $\leq O(\frac{n^{2\gamma+\frac{\tau}{2}(1+\frac{\gamma}{2})}}{n^{a/2}}\cdot\|{\mathbf{V}}\|_{\operatorname{op}})$ (26)
$\|{\mathbf{\widehat{T}}_{\textrm{hyp}}-\mathbf{T}}\|_{\max}$ $\leq O\big{(}\frac{n^{\frac{17\gamma}{3}}(\log n)^{\frac{1}{6}}}{n^{a/6}}\cdot\|{\mathbf{V}}\|_{\operatorname{op}}\big{)}$ (27)
$\|{\mathbf{\widehat{T}}_{\textrm{thin}}-\mathbf{T}}\|_{\max}$ $\leq O\big{(}\frac{n^{2\gamma}\sqrt{d\log(n\|{\mathbf{V}}\|_{\max})}\log n}{n^{a}}\cdot\|{\mathbf{V}}\|_{2,\infty})$ (28)
with $O(dn^{1+a})$ runtime and probability at least $\frac{1}{2}$. The Thinformer guarantee exhibits four improvements over its predecessors. First, it establishes a significantly faster error decay rate ($n^{-a}$ versus $n^{-a/2}$ or $n^{-a/6}$) for a given subquadratic runtime $n^{1+a}$. Second, it reduces the dependence on the error inflation factor $\gamma$. Third, like the HyperAttention guarantee, it eliminates all dependence on the KDEformer penalty parameter $\tau$. Finally, it reduces dependence on the value matrix by a factor of $\frac{\|{\mathbf{V}}\|_{\operatorname{op}}}{\|{\mathbf{V}}\|_{2,\infty}}\in[1,\sqrt{n}]$.
Put otherwise, with bounded $\|{\mathbf{V}}\|_{2,\infty}$, $\mathbf{\widehat{T}}_{\textrm{thin}}$ can provide consistent (i.e., $\|{\mathbf{\widehat{T}}_{\textrm{thin}}-\mathbf{T}}\|_{\max}\to 0$ as $n\to\infty$) subquadratic estimation whenever $\gamma$ is bounded away from $1/2$ and guarantee, for example, $O(\frac{1}{\sqrt{n}})$ error in $\widetilde{O}(dn^{\frac{3}{2}+2\gamma})$ time. In contrast, the $\mathbf{\widehat{T}}_{\textrm{kde}}$ and $\mathbf{\widehat{T}}_{\textrm{hyp}}$ bounds require quadratic runtime to guarantee $O(\frac{1}{\sqrt{n}})$ error in the best case (${\|{\mathbf{V}}\|_{\operatorname{op}}}=O(1)$) and cannot guarantee consistent subquadratic estimation in the worst case (${\|{\mathbf{V}}\|_{\operatorname{op}}}=\Omega(\sqrt{n})$).

### v7

# Low-Rank Thinning

@@ 1 Introduction
This work is about thinning, finding a small set of representative points to accurately summarize a larger dataset. Here, we use the term “dataset” liberally to refer to any collection of points, be they experimental observations (as in Sec. 6), stochastic gradients (as in Sec. 5), or key-value pairs (as in Sec. 4). State-of-the-art thinning techniques provably improve upon uniform subsampling but only for restricted classes of kernel-based quality measures and with pessimistic dependence on the data dimension (see, e.g., Harvey & Samadi, 2014; Phillips & Tai, 2020; Alweiss et al., 2021; Dwivedi & Mackey, 2024, 2022; Shetty et al., 2022; Li et al., 2024). We introduce a new analysis for sub-Gaussian thinning algorithms that applies to any kernel and shows that one can efficiently identify a better-than-uniform set of representative points whenever the kernel or data matrix is nearly low-rank. This opens the door to a variety of impactful applications including approximate dot-product attention in transformers, accelerating model training through gradient reordering, and distinguishing distributions with deep kernels in near-linear time.
In Sec. 2, we introduce our formal definition of thinning, two kernel-based measures of thinning quality, and a suite of candidate thinning algorithms. Sec. 3 presents our main theorem relating thinning quality to low-rank properties of a dataset and its induced kernel matrix. In Sec. 4, we translate the problem of approximating attention into a thinning problem and develop a practical solution, Thinformer, with state-of-the-art quality guarantees. In Sec. 5, we develop a thinned stochastic gradient reordering rule that provably accelerates training and bridges the theory-practice gaps left open by prior work. Finally, in Sec. 6 we derive new and improved power guarantees for practical thinned hypothesis tests that distinguish distributions in near-linear time. Each section also includes its own discussion of related work.
@@ 2 Sub-Gaussian Thinning
We begin with a formal definition of our problem setting. Consider a fixed collection of $n_{\textup{in}}$ input points $\mathcal{X}_{\textup{in}}$ belonging to a potentially larger universe of datapoints $\mathcal{X}\triangleq\{\boldsymbol{x}_{1},\dots,\boldsymbol{x}_{n}\}$. The aim of a thinning algorithm is to select $n_{\textup{out}}$ points from $\mathcal{X}_{\textup{in}}$ that together accurately summarize $\mathcal{X}_{\textup{in}}$. This is formalized by the following definition.
@@
[…]
@@
When $\mathcal{X}\subset\real^{d}$, we use $\mathbf{X}\triangleq[\boldsymbol{x}_{1},\ldots,\boldsymbol{x}_{n}]^{\top}\in\real^{n\times d}$ to denote the input point matrix so that
@@
Given two distributions $\mu,\tilde{\mu}$ and a reproducing kernel $\mathbf{k}$ (Steinwart & Christmann, 2008, Def. 4.18), the associated kernel maximum mean discrepancy (MMD) is the worst-case difference in means for functions in the unit ball $\mathbb{B}_{\mathbf{k}}$ of the associated reproducing kernel Hilbert space:
$\operatorname{MMD}_{\mathbf{k}}(\mu,\tilde{\mu})$ $\triangleq\sup_{f\in\mathbb{B}_{\mathbf{k}}}|\mathbb{E}_{\boldsymbol{x}\sim\mu}f(\boldsymbol{x})-\mathbb{E}_{\boldsymbol{x}\sim\tilde{\mu}}f(\boldsymbol{x})|.$ (3)
When $\mu=\mathbb{P}_{\textup{in}}$ and $\tilde{\mu}=\mathbb{P}_{\textup{out}}$ as in Def. 1 and $\mathbf{K}\triangleq(\mathbf{k}(\boldsymbol{x}_{i},\boldsymbol{x}_{j}))_{i,j=1}^{n}\in\real^{n\times n}$ denotes the induced kernel matrix, then the MMD can be expressed as a Mahalanobis distance between $\boldsymbol{p}_{\textup{in}}$ and $\boldsymbol{p}_{\textup{out}}$:
@@
We write $\textsc{Alg}\in\mathcal{G}_{\nu,\delta}(\mathbf{K})$ and say Alg is $(\mathbf{K},\nu,\delta)$-sub-Gaussian, if Alg is a thinning algorithm, $\mathbf{K}$ is a symmetric positive semidefinite (SPSD) matrix, $\nu>0$, $\delta\in[0,1)$, and there exists an event $\mathcal{E}$ with probability at least $1-\delta/2$ such that, the input and output probability vectors satisfy
@@ 2.1 Examples of sub-Gaussian thinning algorithms
[…]
@@
Fortunately, uniform subsampling is not the only sub-Gaussian thinning algorithm available. For example, the Kernel Halving (KH$(\delta)$) algorithm of Dwivedi & Mackey (2024) provides a substantially smaller sub-Gaussian parameter, $\nu=O({\sqrt{\log(n_{\textup{out}}/\delta)}}{/n_{\textup{out}}})$, in $n_{\textup{in}}^{2}$ time by selecting one out of every two points and biasing selection toward the point that yields smaller approximation error. Similarly, the KH-Compress$(\delta)$ algorithm of Shetty et al. (2022, Ex. 3) delivers $\nu=O({\sqrt{\log(n_{\textup{out}})\log(n_{\textup{out}}/\delta)}}{/n_{\textup{out}}})$ in only $n_{\textup{out}}^{2}$ time by halving and concatenating coresets of geometrically increasing size until the target output size is met. We derive simplified versions of these algorithms with identical sub-Gaussian constants in Secs. B.2 and B.5 and a linear-kernel variant (LKH$(\delta)$) with $n_{\textup{in}}d$ runtime in Sec. B.3. To round out our set of examples, we show in Sec. B.6.1 that two new thinning algorithms based on the Gram-Schmidt walk of Bansal et al. (2018) yield even smaller $\nu$ at the cost of increased runtime. We call these algorithms Gram-Schmidt Thinning (GS-Thin) and GS-Compress.
@@ 3 Low-rank Sub-Gaussian Thinning
Let us unpack the three components of this result. First, Thm. 1 provides a high-probability $O(\nu\sqrt{\log(|\mathcal{I}|)})$ bound 13 on the KMS for any kernel and any sub-Gaussian thinning algorithm on any space. In particular, the non-uniform algorithms of Tab. 1 all enjoy $O(\log(n_{\textup{out}})\sqrt{\log(|\mathcal{I}|)}/n_{\textup{out}})$ KMS, a significant improvement over the $\Omega(1/\sqrt{n_{\textup{out}}})$ KMS of uniform subsampling. The bound 13 follows from the sub-Gaussianity of the thinning algorithm (Def. 3) and the union bound over ±$\pm\boldsymbol{e}_{i}$ for each $i\in\mathcal{I}$.
For datapoints in $\mathbb{R}^{d}$, Thm. 1 also provides a refined $O(\nu\sqrt{\mathrm{rank}({\mathbf{X}_{\mathcal{I}}})\log(R_{\mathcal{I}}L_{\mathbf{K}})})$ bound 15 on KMS. For bounded data, this trades an explicit dependence on the number of query points $|\mathcal{I}|$ for a rank factor that is never larger (and sometimes significantly smaller) than $d$. This refinement follows from a more elaborate chaining argument that frames $(\boldsymbol{e}_{i}^{\top}\mathbf{K}(\boldsymbol{p}_{\textup{in}}-\boldsymbol{p}_{\textup{out}}))_{i\in\mathcal{I}}$ as a sub-Gaussian process (Lem. C.5) and uses the Lipschitzness of $K$ to control its entropy integral. We will make use of these results when approximating dot-product attention in Sec. 4.
Notably, Thm. 3.1 of Phillips & Tai (2020) implies that any thinning algorithm must incur at least $\Omega(\sqrt{d}/n_{\textup{out}})$ KMS error for some dataset in $\mathbb{R}^{d}$ and many common kernels. Meanwhile, our Tab. 1 and Thm. 1 imply that GS-Thin has $\nu=O(1/n_{\textup{out}})$ and hence KMS $O(\sqrt{d}/n_{\textup{out}})$. Taken together, these results imply that no algorithm can have sub-Gaussian constant $\nu=o(1/n_{\textup{out}})$ and that GS-Thin enjoys minimax rate-optimal KMS and a minimax rate-optimal sub-Gaussian constant.
Finally, Thm. 1 provides an $O(\nu\sqrt{r}+\sqrt{\lambda_{r+1}/n_{\textup{out}}})$ high-probability bound on kernel MMD, where the approximate rank parameter $r$ can be freely optimized. We establish this result by projecting $\mathbf{K}^{1/2}(\boldsymbol{p}_{\textup{in}}-\boldsymbol{p}_{\textup{out}})$ onto the first $r$ eigenvectors of $\mathbf{K}$, bounding the residual error in terms of $\lambda_{r+1}$, and bounding the projection magnitude with high probability using the sub-Gaussianity of Alg and a union bound over a finite cover of a Euclidean ball in $\mathbb{R}^{r}$.
When $\mathbf{K}=(\mathbf{k}(\boldsymbol{x}_{i},\boldsymbol{x}_{j}))_{i,j=1}^{n}$ is generated by a finite-rank kernel $\mathbf{k}$, like a linear kernel $\langle{\boldsymbol{x}_{i}},{\boldsymbol{x}_{j}}\rangle$, a polynomial kernel $(1+\langle{\boldsymbol{x}_{i}},{\boldsymbol{x}_{j}}\rangle)^{p}$, or a random Fourier feature kernel (Rahimi & Recht, 2007), this guarantee becomes $O(\nu)$ and improves upon uniform subsampling whenever $\nu=o(1/\sqrt{n_{\textup{out}}})$. In this case, the non-uniform algorithms of Tab. 1 all enjoy $O(\log(n_{\textup{out}})/n_{\textup{out}})$ MMD, a significant improvement over the $\Omega(1/\sqrt{n_{\textup{out}}})$ MMD of uniform subsampling. We will revisit this finite-rank setting when studying stochastic gradient acceleration strategies in Sec. 5.
@@
Altschuler et al. (2019, Thm. 4) additionally showed that Gaussian kernel matrix eigenvalues satisfy
@@
With our core theory in hand, we now turn our attention to a series of impactful applications.
@@ 4 Approximating Attention
We will first use our analysis to accelerate attention approximation in transformers. Dot-product attention lies at the heart of the transformer neural network architecture that has revolutionized natural language processing, computer vision, and speech recognition over the last decade (Vaswani et al., 2017; Dosovitskiy et al., 2021; Dong et al., 2018). Given a collection of query, key, and value vectors $(\boldsymbol{q}_{i},\boldsymbol{k}_{i},\boldsymbol{v}_{i})_{i=1}^{n}$ each in $\mathbb{R}^{d}$, dot-product attention computes the softmax matrix
@@
$\text{for}\ \mathbf{A}_{ij}\triangleq\exp(\frac{\langle{\boldsymbol{q}_{i}},{\boldsymbol{k}_{j}}\rangle}{\sqrt{d}}),\mathbf{D}=\mathop{\mathrm{diag}}(\mathbf{A}\mathbf{1}_{n}),\text{ and }\mathbf{V}_{ij}\triangleq\boldsymbol{v}_{ij}.$ (23)
@@ 4.1 Thinning attention in theory
[…]
@@
Table 3: Quality of T2T-ViT attention approximations on ImageNet. We report mean Top-111 accuracy ±1$\pm$ 1± 1 standard deviation across five random seeds and mean forward pass runtime ±1$\pm$ 1± 1 standard deviation across 505050 batches of 646464 images.
@@
$\|{\mathbf{\widehat{T}}-\mathbf{T}}\|_{\max}\leq\!\frac{c\exp(\frac{2R^{2}}{\sqrt{d}})\|{\mathbf{V}}\|_{2,\infty}\sqrt{\log_{2}(n_{\textup{out}})\log({8n_{\textup{out}}\log_{2}\frac{n_{\textup{in}}}{n_{\textup{out}}}})}}{n_{\textup{out}}}$ (24)
@@
the KDEformer approximation (Zandieh et al., 2023, Cor. 3.6) with $\tau=0.173+o(1)$, the HyperAttention approximation (Han et al., 2024, Thm. 1) with no masking, and the Thinformer approximation (Thm. 2) guarantee the $\|{\mathbf{\widehat{T}}-\mathbf{T}}\|_{\max}$ bounds of Tab. 2 with $O(dn^{1+a})$ runtime and probability at least $\frac{1}{2}$. The Thinformer guarantee exhibits four improvements over its predecessors. First, it establishes a significantly faster error decay rate ($n^{-a}$ versus $n^{-a/2}$ or $n^{-a/6}$) for a given subquadratic runtime $n^{1+a}$. Second, it reduces the dependence on the error inflation factor $\gamma$. Third, like the HyperAttention guarantee, it eliminates all dependence on the KDEformer penalty parameter $\tau$. Finally, it reduces dependence on the value matrix by a factor of $\frac{\|{\mathbf{V}}\|_{\operatorname{op}}}{\|{\mathbf{V}}\|_{2,\infty}}\in[1,\sqrt{n}]$.
Put otherwise, with bounded $\|{\mathbf{V}}\|_{2,\infty}$, $\mathbf{\widehat{T}}_{\textrm{thin}}$ can provide consistent (i.e., $\|{\mathbf{\widehat{T}}_{\textrm{thin}}-\mathbf{T}}\|_{\max}\to 0$ as $n\to\infty$) subquadratic estimation whenever $\gamma$ is bounded away from 1/21/21 / 2 and guarantee, for example, $O(\frac{1}{\sqrt{n}})$ error in $\widetilde{O}(dn^{\frac{3}{2}+2\gamma})$ time. In contrast, the $\mathbf{\widehat{T}}_{\textrm{kde}}$ and $\mathbf{\widehat{T}}_{\textrm{hyp}}$ bounds require quadratic runtime to guarantee $O(\frac{1}{\sqrt{n}})$ error in the best case (${\|{\mathbf{V}}\|_{\operatorname{op}}}=O(1)$) and cannot guarantee consistent subquadratic estimation in the worst case (${\|{\mathbf{V}}\|_{\operatorname{op}}}=\Omega(\sqrt{n})$).
Table 4: Quality of BigGAN attention approximations for image generation. We report Frechet Inception Distance (FID) with the ImageNet validation set, Inception Scores (IS), and mean forward pass runtime ± 1 standard deviation across 10 batches of 32 images. A lower FID or higher IS indicates better image generation quality.
[…]